# ORCID: $0$-0003-2317-2297

### v1

@@ 6.1 Threats to internal validity
    Lastly, the study uses Natural Language Generation (NLG) as a tool for generating natural language summaries. While functional, the generated summaries are relatively primitive in their current form. The quality of these summaries could potentially be enhanced by leveraging more advanced tools or techniques, such as SWUM (Software Word Usage Model) or grammar-based generators, which may offer more sophisticated and contextually accurate outputs. Future work could explore the integration of such tools to improve the overall quality and readability of the generated summaries
@@ 8 Conclusion & Future Works
Appendix A Reproducibility
For the purpose of reproducibility of our results, we have released our complete implementation with the annotated reference set and the result files to the public as an open-source project via our online appendix at https://github.com/najamnazar/DPS-Design-Pattern-Summariser. We have collected all projects in the corpus and have committed them to a GitHub repository which is accessible at https://github.com/SamSike/DPS-Corpus.

### v3

@@ 7.1 Threats to internal validity
    Lastly, the study uses Natural Language Generation (NLG) as a tool for generating natural language summaries. While functional, the generated summaries are relatively primitive in their current form. The quality of these summaries could potentially be enhanced by leveraging more advanced tools or techniques, such as SWUM (Software Word Usage Model) or grammar-based generators, which may offer more sophisticated and contextually accurate outputs. Future work could explore the integration of such tools to improve the overall quality and readability of the generated summaries.
  4. 4.
    Our approach assumes that the design-pattern detections provided to the summariser are correct. Any inaccuracies or false positives in pattern detection may result in incomplete or misleading summaries. To mitigate this threat, we manually verified a sample of the detected pattern instances against reference implementations. Future work will explicitly quantify how detection errors affect summary quality and explore coupling the summarisation pipeline with probabilistic confidence estimates from the detector.
@@ 8.2 Design Pattern Identification
extcolorblueRecent advances in large language models (e.g., GPT-4, CodeLlama) have demonstrated strong general-purpose summarisation capabilities. Applying such models to design-pattern documentation is a promising direction for future work. However, because LLM outputs are sensitive to prompt phrasing, temperature, and model version, they are less reproducible and less interpretable than feature-driven approaches like DPS. Future studies could integrate both paradigms, combining explicit structural features with LLM-based contextual reasoning to balance precision and creativity.
@@ Acknowledgements.
Reproducibility
For the purpose of reproducibility of our results, we have released our complete implementation with the annotated reference set and the result files to the public as an open-source project via our online appendix at https://github.com/najamnazar/designpatternsummariser. We have collected all projects in the corpus and have committed them to a GitHub repository which is accessible at https://github.com/SamSike/DPS-Corpus.
Appendix - Declaration
Funding
This project was not supported by any internal or external funding.
Ethical Approval
Not Applicable
Informed Consent
All participants provided informed consent before taking part in this research.
Author Contributions
Najam Nazar was responsible for data collection, data analysis, experimentation (including code development), and manuscript writing. Sameer Sikka contributed to the experimentation and corrected bugs and errors. Christoph Treude was involved in manuscript writing and review.
Data Availability Statement
Not Applicable
Conflict of Interest
The authors have no conflicts of interest to declare relevant to this paper’s content.
Clinical trial number
Not Applicable